Given f(x)=4x^2+3x-1 What are the zeroes in the function
I Only got as fas as 0=4(x^2+3x-1)

Answer by Alan3354(69443)

Given f(x)=4x^2+3x-1 What are the zeroes in the function
I Only got as fas as 0=4(x^2+3x-1)
------------
0=4(x^2+3x-1) is not right. You didn't divide all 3 terms by 4.
4x^2+3x-1 = 0
(4x - 1)*(x + 1) = 0
x = -1
x = 1/4